OpenStudy (anonymous):

Write an example of a system of three equations in three variables that has (­3, 5, 2) as a solution. Show that the ordered triple satisfies all three equations.

OpenStudy (solomonzelman):

You can do it backwards :)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle 2(3)-(5)+2(2)=5 }\) \(\color{#000000}{ \displaystyle (3)+(5)-(2)=6 }\) \(\color{#000000}{ \displaystyle 6(3)-2(5)-3(2)=2 }\) \(\color{#000000}{ \displaystyle 2x-y+2z=5 }\) \(\color{#000000}{ \displaystyle x+y-z=6 }\) \(\color{#000000}{ \displaystyle 6x-2y-3z=2 }\) and clearly, \(\color{#000000}{ \displaystyle (x,y,z)=(3,5,2) }\) is a solution.

OpenStudy (solomonzelman):

See what I am doing?

OpenStudy (anonymous):

Ah yes! I had come up with something different (with the solutions 3, 5, and 2 but the results just being 0) but this makes a lot more sense. Thank you!

OpenStudy (solomonzelman):

Yeah, when you are asked to come up with a system that has a particular solution, you can always go becwards like this.

OpenStudy (solomonzelman):

(( You can also tell that there is an infinite number of possible sustems that would satisfy a particular solution. ))

OpenStudy (solomonzelman):

In any case, you welcome:)